[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Apr 24, 2017
Matthew Sardi - NOAA Federal
matthew.sardi at noaa.gov
Mon Apr 24 16:24:57 UTC 2017
NCO will be performing upgrades on the core Internet routers in College
Park, MD on Wednesday, April 26, 2017. Please be prepared for intermittent
Internet and NCEP network outages, and associated effects on systems
throughout the maintenance window.
Start Date-Time: 4/26/2017 1015 UTC (6:15 am ET)
Stop Date-Time: 4/26/2017 2015 UTC (4:15 pm ET)
The period of definite impact to users will be 1015 UTC (6:15 am ET) to
1120 UTC (7:20 am ET). We do not expect a complete outage for the full 50
minutes, but rather intermittent outages of various services during this
The intermittent impacts during the 1015 to 1120 UTC period will affect
Remote NCEP Centers, including NHC:
· Loss of internet service
· Interruption to NCEP VPN access and WCOSS
· Delay of NCEP-hosted email services, such as listservs
There will be a second period of potential impact from 1900 UTC (3pm ET) to
2015 UTC (4:15 pm ET). The work during this 75 minutes should not result
in any outages, but it is a high risk period of work that we want users to
be aware of.
NHC/Technology & Science Branch
305-229-4496 <(305)%20229-4496> (Desk)
631-833-2965 <(631)%20833-2965> (Cell)
On Mon, Apr 24, 2017 at 12:09 PM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
> Here is today's summary of NHC computer operations:
> --- Matt Sardi
> 1. TSB opened an 'Urgent' helpdesk ticket on TAFB surface desk workstation
> CUMULUS, which continues to suffer serious performance degradation. NCO
> will be making some system changes and rebooting CUMULUS this afternoon.
> 1. WFO storm surge testing will occur this week (Mon-Thu). Mike Brennan
> will submit a test psurge run each day, and then the WFOs will make sure
> they can ingest it. Chris Mello will conduct daily conference calls with
> the WFOs.
> *Other items:*
> 1. Mark DeMaria and Jamie Rhome are attend the Coastal Inundation Summit
> in Washington, D.C. this week. Craig Mattocks is acting TSB Chief.
> 2.GOES-16 'Mode 4' (continuous full disk at 5 minutes, no conus/mesos)
> testing ended at 12Z today. However, Mike Brennan reports that no new
> images have ingested since ~00Z on 4/22. TSB will investigate.
> 3. NCO will be activating the RHEL6 replacement for CARCAH's 'scout2'
> workstation today or tomorrow, so Steve Feuer and Warren Madden can fully
> test functionality.
> *NHC RFCs:*
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts